The sum of two numbers is 20. The difference of the two numbers is twice the smaller number. Write a system of equations that will determine the two numbers.
Answer by checkley75(3666)

X+Y=20 THIS IS THE SUM OF THE TWO NUMBERS WITH X BEING THE SMALLER NUMBER.
Y-X=2X THIS IS THE DIFFERENCE THAT IS TWICE THE SMALLER NUMBER.
YOU DON'T ASK TO SOLVE FOR THE TWO NUMBERS BUT I WILL SOLVE THEM FOR YOU.
X+Y=20 OR X=20-Y NOW SUBSTITUTE (20-Y) FOR X IN THE OTHER EQUATION THUS:
Y-(20-Y)=2(20-Y)
Y-20+Y=40-2Y
2Y+2Y=40+20
4Y=60
Y=60/4
Y=15 ANSWER.
X+15=20
X=20-15
X=5 ANSWER.
PROOF
15-5=2*5
10=10